2 What is the distance traveled by the car in Problem # 1, in 3 seconds?

2

What is the distance traveled by the car in Problem # 1, in 3 seconds?

Grade:12

1 Answers

Yash Jain
9 years ago
s = ut + ½*at2 (2nd eqn of motion)
In this case, u = 0m/s, a = 4m/s2, t = 3s
Therefore, s = 0 + ½*4*32 = 18m